3 over x +6 = negative 3\4
Answer by unlockmath(1688)

Hello,
Let's write this like this:
3/x +6 = -3\4 Now, let's get rid of the fractions by multiplying both sides by 4x
12+24x=-3x Now Subtract 12 and add 3x to both sides to give us:
27x=-12 Divide both sides by 27 will be:
x=-12/27 or rewritten as -4/9
Plug this answer into the original problem to see if it works.
Make sense?
RJ
Check out a new book at:
www.math-unlock.com